OpenStudy (anonymous):

Which is the slope of the following equation ? y = -x - 2

OpenStudy (ddcamp):

y=mx+b m: The slope b: The y-intercept What's m in your equation?

OpenStudy (anonymous):

thats how the question was written no slope(m) included

OpenStudy (anonymous):

When there is no specific number in front of a variable, it is 1.

OpenStudy (ddcamp):

There is a - sign in front of the x, which is the same as saying (-1)x

OpenStudy (anonymous):

What is the slope of the line that passes through (3, 5) and (6, 10)?

OpenStudy (anonymous):

would it be \[\frac{ 3 }{ 5 }? \]

OpenStudy (ddcamp):

Close, but the slope is rise/run, and you have run/rise.

OpenStudy (anonymous):

\[\frac{ 5 }{ 3 }\]